OpenStudy (anonymous):

evaluate analytically : \frac{ sqrt{5x+5h} - sqrt{5x} }{ h }

OpenStudy (campbell_st):

I'll assume its 1st principals... with a limit you need to multiply by 1 in a different form \[\lim_{h \rightarrow 0} \frac{\sqrt{x + h} - \sqrt{x}}{h} \times \frac{\sqrt{x + h}+\sqrt{x}}{\sqrt{x + h} + \sqrt{x}}\] which becomes \[\lim_{h \rightarrow 0} \frac{x + h - x}{h(\sqrt{x + h} - \sqrt{x})}\] which should be easy to finish...

OpenStudy (anonymous):

i did and got the wrong answer.. idk where i went wrong

OpenStudy (campbell_st):

ok... so it becomes \[\lim_{h \rightarrow 0} \frac{h}{h(\sqrt{x + h} + \sqrt{x})}\] cancel the h's from the numerator and denominator gives \[\lim_{h \rightarrow 0} \frac{1}{\sqrt{x +h} + \sqrt{x}}\] now apply the limit for h... subsitute h = 0 and what do you get

OpenStudy (anonymous):

so sorry for the late response.. i didnt think you answered.. but thank you for the help